OpenStudy (saranya):

Can anyone help out this question....? Arrange the following functions in increasing order of growth rate. ie. if g(n) follows f(n) in your list, then f(n) is necessarily O(g(n)). a)2^2n b)2^n^2 c)n^2log(n) d)n e)n^2n

ganeshie8 (ganeshie8):

put n=10000 and see

OpenStudy (anonymous):

i am confused by the last part

OpenStudy (saranya):

increasing order : d>a>b>c>e is this correct ?

OpenStudy (anonymous):

some should be easy \(f(n)=n\) is first on the list

OpenStudy (anonymous):

i think you have it right, but i would have written the inequalities the other way

OpenStudy (anonymous):

hmm maybe i am wrong

OpenStudy (anonymous):

looks like \(2^{x^2}\) grows faster than \(x^{2x}\)

OpenStudy (saranya):

can yu give me ur ans. i will work out n c to t :)

OpenStudy (saranya):

give ur ans in increasing order :)

OpenStudy (anonymous):

i didn't do any work, i just checked with wolfram if \(\lim_{x\to\infty}\frac{2^{x^2}}{x^{2x}}=\infty\) then \(2^{x^2}\) grows faster

OpenStudy (saranya):

yup so thats what i mentioned ryt ? :)

OpenStudy (anonymous):

a)2^2n b)2^n^2 c)n^2log(n) d)n e)n^2n least is d then looks like c from the last wolfram link

OpenStudy (anonymous):

then my guess is a then e then b but i was cheating and going fast check each gotta run

OpenStudy (saranya):

okay :)

OpenStudy (saranya):

yu were correct dude :)